# Impact of flavour coupling on $SO(10)$-inspired leptogenesis

### v1

@@ 2.1 Seesaw mechanism and low energy neutrino data
[…]
where $D_{m_{\ell}}\equiv{\rm diag}(m_{e},m_{\mu},m_{\tau})$ is the diagonal charged lepton mass matrix, $D_{M}\equiv{\rm diag}(M_{1},M_{2},M_{3})$ is the diagonal Majorana mass matrix and $m_{D}=hv$ is the neutrino Dirac mass matrix. In the seesaw limit, for $M\gg m_{D}$, the mass spectrum splits into two sets of Majorana neutrino eigenstates, a light set, $\nu_{1}$, $\nu_{2}$ and $\nu_{3}$, with masses $m_{1}\leq m_{2}\leq m_{3}$, given by the seesaw formula [2]
$D_{m}=U^{\dagger}\,m_{D}\,{1\over D_{M}}\,m_{D}^{T}\,U^{\star}\,,$ (4)
with $D_{m}={\rm diag}(m_{1},m_{2},m_{3})$, and a heavy set, $N_{1}$, $N_{2}$ and $N_{3}$, with masses almost coinciding with the three $M_{I}$ in $D_{M}$. The matrix $U$, diagonalising the light neutrino mass matrix $m_{\nu}=-m_{D}\,M^{-1}\,m_{D}^{T}$ in the weak basis, can be identified with the PMNS lepton mixing matrix. This can be parameterised in terms of the usual three mixing angles $\theta_{ij}$, the Dirac phase $\delta$ and the Majorana phases $\rho$ and $\sigma$, as
$U=\left(\begin{array}[]{ccc}c_{12}\,c_{13}&s_{12}\,c_{13}&s_{13}\,e^{-{\rm i}\,\delta}\\ -s_{12}\,c_{23}-c_{12}\,s_{23}\,s_{13}\,e^{{\rm i}\,\delta}&c_{12}\,c_{23}-s_{12}\,s_{23}\,s_{13}\,e^{{\rm i}\,\delta}&s_{23}\,c_{13}\\ s_{12}\,s_{23}-c_{12}\,c_{23}\,s_{13}\,e^{{\rm i}\,\delta}&-c_{12}\,s_{23}-s_{12}\,c_{23}\,s_{13}\,e^{{\rm i}\,\delta}&c_{23}\,c_{13}\end{array}\right)\,{\rm diag}\left(e^{i\,\rho},1,e^{i\,\sigma}\right)\,,$ (5)
where $s_{ij}\equiv\sin\theta_{ij}$ and $c_{ij}\equiv\cos\theta_{ij}$. As we said, successful SO10INLEP cannot be realised within IO, compatibly with the latest upper bounds on neutrino masses and atmospheric neutrino mixing angle. For this reason we only consider NO. In this case latest global analyses of neutrino oscillation experiment results, including atmospheric neutrino data from Super-Kamiokande and IceCube collaborations, find for the mixing angles and leptonic Dirac phase $\delta$ the following best fit values, $1\sigma$ errors and $3\sigma$ intervals [23]:
@@
No neutrinoless double beta ($0\nu\beta\beta$) decay signal has been detected so far. This implies that there are no experimental constraints on the Majorana phases and that experiments can only place an upper bound on the effective $0\nu\beta\beta$ neutrino mass $m_{ee}\equiv|m_{\nu ee}|$. The most stringent one has been set by the KamLAND-Zen collaboration, that found [42]
@@
where the range accounts for nuclear matrix element uncertainties. This translates into an upper bound on the lightest neutrino mass $m_{1}\leq(84\mbox{--}353)\,{\rm meV}$ (90%  C.L.). When this is compared with the upper bound Eq. (2) from cosmological observations (assuming $\Lambda$CDM), one can clearly see how the latter is much more stringent. [^4] It is even more stringent than it looks like, considering that the cosmological upper bound is at $95\%$ C.L. and the upper bound from $0\nu\beta\beta$ at $90\%$ C.L. . However, a future $0\nu\beta\beta$ positive signal with a measurement of $m_{ee}$ would not just provide information on the absolute neutrino mass scale but also some partial information on the Majorana phases.
@@
on the effective electron neutrino mass [43]. Since this falls in the quasi-degenerate limit, it translates into an equal upper bound on $m_{1}$.
@@ 2.2 Combining seesaw and $SO(10)$-inspired conditions
Let us now briefly review seesaw models when $SO(10)$-inspired conditions [9] are imposed on the neutrino Dirac mass matrix. This can be diagonalised with the so-called singular value decomposition (sometimes also referred to as bi-unitary parameterisation) as
@@
Parameterising the neutrino Dirac masses $m_{Di}$ in terms of the up quark masses, [^5] For the values of the up-quark masses at the scale of leptogenesis ($\sim 10^{11}\,{\rm GeV}$), we adopt $(m_{\rm up},m_{\rm charm},m_{\rm top})=(1\,{\rm MeV},400\,{\rm MeV},100\,{\rm GeV})$ [44].
@@
we impose $SO(10)$-inspired conditions [9, 10, 16] defined as
[…]
With the latter we imply that using for the matrix $V_{L}$ the same parameterisation as for the leptonic mixing matrix $U$, the three mixing angles $\theta_{12}^{L}$, $\theta_{23}^{L}$ and $\theta_{13}^{L}$ do not have values much larger than the three mixing angles in the CKM matrix and in particular $\theta_{12}^{L}\lesssim\theta_{c}\simeq 13^{\circ}$, where $\theta_{c}$ is the Cabibbo angle.
@@
From the analytical procedure discussed in [10, 12, 33], one finds simple expressions for the three RH neutrino masses,
@@
with the three phases in $D_{\phi}\equiv{\rm diag}(e^{-i\,{\Phi_{1}\over 2}},e^{-i{\Phi_{2}\over 2}},e^{-i{\Phi_{3}\over 2}})$ given by [33]
@@
One can also derive an expression for the orthogonal matrix. Starting from its definition $\Omega=D_{m}^{-{1\over 2}}\,U^{\dagger}\,m_{D}\,D_{M}^{-{1\over 2}}$ [45] that, using Eq. (12), becomes [16]
$\Omega=D_{m}^{-{1\over 2}}\,U^{\dagger}\,V_{L}^{\dagger}\,D_{m_{D}}\,U_{R}\,D_{M}^{-{1\over 2}}\,$ (21)
or, in terms of its matrix elements ($\alpha=e,\mu,\tau$; $i=1,2,3$; $I=1,2,3$), [^6] The different labels denote the three different sets of lepton flavours. While the charged lepton neutrino flavours and the neutrino mass eigenstates form two orthonormal bases, the heavy neutrino lepton flavours cannot in order for the $C\!P$ asymmetries not to all vanish and have successful leptogenesis.
$\Omega_{iJ}\simeq{1\over\sqrt{m_{i}\,M_{J}}}\,\sum_{l=1}^{3}\,m_{Dl}\,U^{\star}_{\alpha i}\,V^{\star}_{L\,l\alpha}\,U_{R\,lJ}\,,$ (22)
one finds [33]
$\Omega\simeq\left(\begin{array}[]{ccc}i\,{(\widetilde{m}_{\nu}\,W^{\star})_{11}\over\sqrt{m_{1}\,\widetilde{m}_{\nu 11}}}&\sqrt{m_{2}\,m_{3}\,(\widetilde{m}_{\nu}^{-1})_{33}\over\widetilde{m}_{\nu 11}}\,\left(W^{\star}_{21}-W^{\star}_{31}\,{{(\widetilde{m}_{\nu}^{-1})_{23}}\over(\widetilde{m}_{\nu}^{-1})_{33}}\right)&{W^{\star}_{31}\over\sqrt{m_{1}\,(\widetilde{m}_{\nu}^{-1})_{33}}}\\ i\,{(\widetilde{m}_{\nu}\,W^{\star})_{12}\over\sqrt{m_{2}\,\widetilde{m}_{\nu 11}}}&\sqrt{m_{1}\,m_{3}\,(\widetilde{m}_{\nu}^{-1})_{33}\over\widetilde{m}_{\nu 11}}\,\left(W^{\star}_{22}-W^{\star}_{32}\,{{(\widetilde{m}_{\nu}^{-1})_{23}}\over(\widetilde{m}_{\nu}^{-1})_{33}}\right)&{W^{\star}_{32}\over\sqrt{m_{2}\,(\widetilde{m}_{\nu}^{-1})_{33}}}\\ i\,{(\widetilde{m}_{\nu}\,W^{\star})_{13}\over\sqrt{m_{3}\,\widetilde{m}_{\nu 11}}}&\sqrt{m_{1}\,m_{2}\,(\widetilde{m}_{\nu}^{-1})_{33}\over\widetilde{m}_{\nu 11}}\,\left(W^{\star}_{23}-W^{\star}_{33}\,{{(\widetilde{m}_{\nu}^{-1})_{23}}\over(\widetilde{m}_{\nu}^{-1})_{33}}\right)&{W^{\star}_{33}\over\sqrt{m_{3}\,(\widetilde{m}_{\nu}^{-1})_{33}}}\end{array}\right)\,,$ (23)
where we defined $W\equiv V_{L}\,U$. The validity of these expressions clearly breaks down in the close vicinity of the two level crossings where either $(\widetilde{m}_{\nu}^{-1})_{33}$, or $\widetilde{m}_{\nu 11}$, or both, vanish. However, as discussed in [12, 33], the first two cases cannot yield successful leptogenesis. The compact spectrum scenario discussed in the introduction correspond to a situation close to the third case. In all three cases the orthogonal matrix entries diverge, a clear indication of the fine tuning that is needed in the the seesaw formula to satisfy neutrino oscillation experimental data.
@@ 2.3 $SO(10)$-inspired leptogenesis
Let us now discuss how the matter-antimatter asymmetry of the universe can be calculated within $SO(10)$-inspired leptogenesis. This can be expressed in terms of the baryon-to-photon number ratio, whose measured value from Planck data (including lensing) combined with external data sets [24], is found
@@
However, some external mechanism might have generated a large value of the initial pre-existing asymmetry, $N_{B-L}^{\rm p,i}$, between the end of inflation and the onset of leptogenesis. In the absence of any washout, this would translate into a sizeable value of $\eta_{B}^{\rm p}$ comparable or greater than $\eta_{B}^{\rm exp}$. In this case, then one should also add a strong thermal leptogenesis condition, requiring that the initial pre-existing asymmetry is efficiently washed out by seesaw neutrino inverse processes. In Section 4 we discuss how a subset of the solutions satisfying successful SO10INLEP can also simultaneously satisfy this strong thermal leptogenesis condition realising ST-SO10INLEP [30]. We also show how this subset is not disrupted by the inclusion of flavour coupling but just slightly modified. For the time being, we will simply assume that the initial pre-existing asymmetry contribution is absent or in any case negligible.
The contribution to $\eta_{B}$ from leptogenesis, can be calculated as [46]
@@
This expression accounts for sphaleron conversion [47, 48, 49, 50] and photon dilution. The second numerical expression holds when the abundances $N_{X}$ are normalised in a portion of comoving volume such that the ultra-relativistic thermal equilibrium abundance of a RH neutrino $N_{N_{i}}^{\rm eq}(T\gg M_{i})=1$. Successful leptogenesis requires $\eta_{B}^{\rm lep}=\eta_{B}^{\rm exp}$, where $\eta_{B}^{\rm exp}$ is the measured value given in Eq. (24).
The final $B-L$ asymmetry from leptogenesis can be calculated, neglecting flavour coupling, as the sum of the three contributions, one from each (charged lepton) flavour, explicitly [19, 51, 36, 52]:
@@
[…]
@@
are the $N_{2}$-flavoured $C\!P$ asymmetries, with $\Gamma_{2}\equiv\sum_{\alpha}\Gamma_{2\alpha}$ and $\overline{\Gamma}_{2}\equiv\sum_{\alpha}\,\overline{\Gamma}_{2\alpha}$. Finally, we defined $K_{2\tau_{2}^{\bot}}\equiv K_{2e}+K_{2\mu}$ and $\varepsilon_{2\tau_{2}^{\bot}}\equiv\varepsilon_{2e}+\varepsilon_{2\mu}$.
Using the bi-unitary parameterisation Eq. (12) for $m_{D}$, the flavoured decay parameters can be expressed as [^8] One can also express the decay parameters in terms of the orthogonal matrix: $K_{I\alpha}=\left|\sum_{j}\,\sqrt{m_{j}\over m_{\star}}\,U_{\alpha j}\,\Omega_{jI}\right|^{2}\,,\hskip 14.22636pt\mbox{\rm and}\hskip 14.22636ptK_{I}=\sum_{I}\,{m_{j}\over m_{\star}}\,|\Omega_{jI}|^{2}\,.$ (30) However, in SO10INLEP, the bi-unitary parameterisation clearly provides the primary way to calculate all leptogenesis parameters, the $K_{I\alpha}$’s and the $\varepsilon_{2\alpha}$’s, since, as one can see from Eqs. (16) and (17), one gets directly RH neutrino masses $M_{I}$ and RH neutrino mixing matrix $U_{R}$ in terms of low energy neutrino parameters in $m_{\nu}$, neutrino Dirac masses $m_{Di}$ and parameters in $V_{L}$.
@@
The flavoured $C\!P$ asymmetries can be calculated using [53]
@@
with $\widetilde{m}_{2}\equiv(m_{D}^{{\dagger}}\,m_{D})_{22}/M_{2}$. However, the first term is important only when RH neutrino masses are quasi-degenerate, close to crossing level solutions. In our case such solutions are falling in region of parameter spaces that are now excluded by experiments [^9] For example, a crossing level solution where $\widetilde{m_{\nu}}^{-1}_{33}=0$ appears at $m_{1}\simeq 0.1\,{\rm eV}$ for the range of values of $\theta_{23}$ we are considering. In any case, in the code we use the full expression for $\varepsilon_{2\alpha}$, also including the contribution from the self-energy diagram to the $C\!P$ asymmetries. and one can reliably use the approximate expression [33]
@@
Notice that in general the asymmetry within minimal leptogenesis with three RH neutrinos would depend on 18 quantities: $\varepsilon_{I\alpha},K_{I\alpha}$ ($I=1,2,3$ and $\alpha=e,\mu,\tau$). However, from Eq. (27) one can see that in SO10INLEP, since it realises a $N_{2}$-leptogenesis scenario, the asymmetry depends only on nine quantities, $\varepsilon_{2\alpha}\,K_{2\alpha},K_{1\alpha}\,(\alpha=e,\mu,\tau)$, while there is no dependence on $\varepsilon_{1\alpha},\varepsilon_{3\alpha},K_{3\alpha}(\alpha=e,\mu,\tau)$. These nine quantities can be expressed in terms of the seesaw parameters. As we have seen, using the bi-unitary parameterisation and the seesaw formula, one obtains Eqs. (16) and (17) for the three Majorana masses and the RH neutrino mixing matrix $U_{R}$ in terms of the six parameters in $V_{L}$, the nine parameters in the low energy neutrino matrix $m_{\nu}$ and the three Dirac neutrino masses $m_{Di}$. Therefore, in this way, the three RH neutrino masses and the six parameters in $U_{R}$ are traded off with the 9 low energy neutrino parameters. Moreover, the $\varepsilon_{2\alpha}$’s depend, approximately, only on $m_{D2}$ and the flavoured decay parameters do not depend on the Dirac neutrino masses $m_{D1},m_{D2}$ and $m_{D3}$. In this way, ultimately, one obtains for the final baryon-to-photon number ratio an expression of the form

### v2

@@ 2.1 Seesaw mechanism and low energy neutrino data
${\cal L}^{m}_{{\ell}+\nu}=\,\overline{\alpha_{L}}\,D_{m_{\ell}}\,\alpha_{R}+\overline{\nu_{\alpha L}}\,m_{D\alpha I}\,N_{RI}+{1\over 2}\,\overline{N^{c}_{RI}}\,D_{M}\,N_{RI}+\mbox{\rm h.c.}\,$, (3)
where $D_{m_{\ell}}\equiv{\rm diag}(m_{e},m_{\mu},m_{\tau})$ is the diagonal charged lepton mass matrix, $D_{M}\equiv{\rm diag}(M_{1},M_{2},M_{3})$ is the diagonal Majorana mass matrix and $m_{D}=hv$ is the neutrino Dirac mass matrix. In the seesaw limit, for $M\gg m_{D}$, the mass spectrum splits into two sets of Majorana neutrino eigenstates, a light set, $\nu_{1}$, $\nu_{2}$ and $\nu_{3}$, with masses $m_{1}\leq m_{2}\leq m_{3}$, given by the seesaw formula seesaw
$D_{m}=U_{\nu}^{\dagger}\,m_{D}\,{1\over D_{M}}\,m_{D}^{T}\,U_{\nu}^{\star}\,$, (4)
with $D_{m}={\rm diag}(m_{1},m_{2},m_{3})$, and a heavy set, $N_{1}$, $N_{2}$ and $N_{3}$, with masses almost coinciding with the three $M_{I}$ in $D_{M}$. The matrix $U_{\nu}$, diagonalising the light neutrino mass matrix $m_{\nu}=-m_{D}\,M^{-1}\,m_{D}^{T}$ in the weak basis, can be identified with the PMNS lepton mixing matrix. This can be parameterised in terms of the usual three mixing angles $\theta_{ij}$, the Dirac phase $\delta$ and the Majorana phases $\rho$ and $\sigma$, as
$U_{\nu}=\left(\begin{array}[]{ccc}c_{12}\,c_{13}&s_{12}\,c_{13}&s_{13}\,e^{-{\rm i}\,\delta}\\ -s_{12}\,c_{23}-c_{12}\,s_{23}\,s_{13}\,e^{{\rm i}\,\delta}&c_{12}\,c_{23}-s_{12}\,s_{23}\,s_{13}\,e^{{\rm i}\,\delta}&s_{23}\,c_{13}\\ s_{12}\,s_{23}-c_{12}\,c_{23}\,s_{13}\,e^{{\rm i}\,\delta}&-c_{12}\,s_{23}-s_{12}\,c_{23}\,s_{13}\,e^{{\rm i}\,\delta}&c_{23}\,c_{13}\end{array}\right)\,{\rm diag}\left(e^{i\,\rho},1,e^{i\,\sigma}\right)\,$, (5)
where $s_{ij}\equiv\sin\theta_{ij}$ and $c_{ij}\equiv\cos\theta_{ij}$. As we said, successful SO10INLEP cannot be realised within IO, compatibly with the latest upper bounds on neutrino masses and atmospheric neutrino mixing angle. For this reason we only consider NO. In this case latest global analyses of neutrino oscillation experiment results, including atmospheric neutrino data from Super-Kamiokande and IceCube collaborations, find for the mixing angles and leptonic Dirac phase $\delta$ the following best fit values, $1\sigma$ errors and $3\sigma$ intervals nufit24 :
@@
No neutrinoless double beta ($0\nu\beta\beta$) decay signal has been detected so far. This implies that there are no experimental constraints on the Majorana phases and that experiments can only place an upper bound on the effective $0\nu\beta\beta$ neutrino mass $m_{ee}\equiv|m_{\nu ee}|$. The most stringent one has been set by the KamLAND-Zen collaboration, that found KamLAND-Zen:2024eml
@@
[…]
@@
on the effective electron neutrino mass Katrin:2024tvg . Since this falls in the quasi-degenerate limit, it translates into an equal upper bound on $m_{1}$.
@@ 2.2 Combining seesaw and $SO(10)$-inspired conditions
Let us now briefly review seesaw models when $SO(10)$-inspired conditions SO10inspired are imposed on the neutrino Dirac mass matrix. This can be diagonalised with the so-called singular value decomposition (sometimes also referred to as bi-unitary parameterisation) as
@@
Parameterising the neutrino Dirac masses $m_{Di}$ in terms of the up quark masses, [^5] For the values of the up-quark masses at the scale of leptogenesis ($\sim 10^{11}\,{\rm GeV}$), we adopt $(m_{\rm up},m_{\rm charm},m_{\rm top})=(1\,{\rm MeV},400\,{\rm MeV},100\,{\rm GeV})$ fusaokakoide .
@@
we impose $SO(10)$-inspired conditions SO10inspired , Akhmedov:2003dg , riotto1 defined as
$\mbox{\rm(i)}\;\alpha_{i}={\cal O}(1)\,;\;\;\;\;\;\;\;\;\;\;\;\;\;\mbox{\rm(ii)}\;\;I\leq V_{L}\lesssim V_{CKM}\,.$ (14)
With the latter we imply that using for the matrix $V_{L}$ the same parameterisation as for the leptonic mixing matrix $U_{\nu}$, the three mixing angles $\theta_{12}^{L}$, $\theta_{23}^{L}$ and $\theta_{13}^{L}$ do not have values much larger than the three mixing angles in the CKM matrix and in particular $\theta_{12}^{L}\lesssim\theta_{c}\simeq 13^{\circ}$, where $\theta_{c}$ is the Cabibbo angle.
@@
From the analytical procedure discussed in Akhmedov:2003dg , decrypting , full , one finds simple expressions for the three RH neutrino masses,
@@
with the three phases in $D_{\phi}\equiv{\rm diag}(e^{-i\,{\Phi_{1}\over 2}},e^{-i{\Phi_{2}\over 2}},e^{-i{\Phi_{3}\over 2}})$ given by full
@@
One can also derive an expression for the orthogonal matrix. Starting from its definition $\Omega=D_{m}^{-{1\over 2}}\,U_{\nu}^{\dagger}\,m_{D}\,D_{M}^{-{1\over 2}}$ Casas:2001sr that, using Eq. (12), becomes riotto1
$\Omega=D_{m}^{-{1\over 2}}\,U_{\nu}^{\dagger}\,V_{L}^{\dagger}\,D_{m_{D}}\,U_{R}\,D_{M}^{-{1\over 2}}\,$ (21)
or, in terms of its matrix elements ($\alpha=e,\mu,\tau$; $i=1,2,3$; I=1,2,3I=1,2,3), [^6] The different labels denote the three different sets of lepton flavours. While the charged lepton neutrino flavours and the neutrino mass eigenstates form two orthonormal bases, the heavy neutrino lepton flavours cannot in order for the CPC\!P asymmetries not to all vanish and have successful leptogenesis.
$\Omega_{iJ}\simeq{1\over\sqrt{m_{i}\,M_{J}}}\,\sum_{l=1}^{3}\,m_{Dl}\,U^{\star}_{\nu\,\alpha i}\,V^{\star}_{L\,l\alpha}\,U_{R\,lJ}\,$, (22)
one finds full
$\hskip-25.60747pt\Omega\simeq\left(\begin{array}[]{ccc}i\,{(\widetilde{m}_{\nu}\,W^{\star})_{11}\over\sqrt{m_{1}\,\widetilde{m}_{\nu 11}}}&\sqrt{m_{2}\,m_{3}\,(\widetilde{m}_{\nu}^{-1})_{33}\over\widetilde{m}_{\nu 11}}\,\left(W^{\star}_{21}-W^{\star}_{31}\,{{(\widetilde{m}_{\nu}^{-1})_{23}}\over(\widetilde{m}_{\nu}^{-1})_{33}}\right)&{W^{\star}_{31}\over\sqrt{m_{1}\,(\widetilde{m}_{\nu}^{-1})_{33}}}\\ i\,{(\widetilde{m}_{\nu}\,W^{\star})_{12}\over\sqrt{m_{2}\,\widetilde{m}_{\nu 11}}}&\sqrt{m_{1}\,m_{3}\,(\widetilde{m}_{\nu}^{-1})_{33}\over\widetilde{m}_{\nu 11}}\,\left(W^{\star}_{22}-W^{\star}_{32}\,{{(\widetilde{m}_{\nu}^{-1})_{23}}\over(\widetilde{m}_{\nu}^{-1})_{33}}\right)&{W^{\star}_{32}\over\sqrt{m_{2}\,(\widetilde{m}_{\nu}^{-1})_{33}}}\\ i\,{(\widetilde{m}_{\nu}\,W^{\star})_{13}\over\sqrt{m_{3}\,\widetilde{m}_{\nu 11}}}&\sqrt{m_{1}\,m_{2}\,(\widetilde{m}_{\nu}^{-1})_{33}\over\widetilde{m}_{\nu 11}}\,\left(W^{\star}_{23}-W^{\star}_{33}\,{{(\widetilde{m}_{\nu}^{-1})_{23}}\over(\widetilde{m}_{\nu}^{-1})_{33}}\right)&{W^{\star}_{33}\over\sqrt{m_{3}\,(\widetilde{m}_{\nu}^{-1})_{33}}}\end{array}\right)\,$, (23)
where we defined $W\equiv V_{L}\,U_{\nu}$. The validity of these expressions clearly breaks down in the close vicinity of the two level crossings where either $(\widetilde{m}_{\nu}^{-1})_{33}$, or $\widetilde{m}_{\nu 11}$, or both, vanish. However, as discussed in decrypting , full , the first two cases cannot yield successful leptogenesis. The compact spectrum scenario discussed in the introduction correspond to a situation close to the third case. In all three cases the orthogonal matrix entries diverge, a clear indication of the fine tuning that is needed in the the seesaw formula to satisfy neutrino oscillation experimental data.
@@ 2.3 $SO(10)$-inspired leptogenesis
Let us now discuss how the matter-antimatter asymmetry of the universe can be calculated within $SO(10)$-inspired leptogenesis. This can be expressed in terms of the baryon-to-photon number ratio, whose measured value from Planck data (including lensing) combined with external data sets Planck:2018vyg , is found
@@
However, some external mechanism might have generated a large value of the initial pre-existing asymmetry, $N_{B-L}^{\rm p,i}$, between the end of inflation and the onset of leptogenesis. In the absence of any washout, this would translate into a sizeable value of $\eta_{B}^{\rm p}$ comparable or greater than $\eta_{B}^{\rm exp}$. In this case, then one should also add a strong thermal leptogenesis condition, requiring that the initial pre-existing asymmetry is efficiently washed out by seesaw neutrino inverse processes. In Section 4 we discuss how a subset of the solutions satisfying successful SO10INLEP can also simultaneously satisfy this strong thermal leptogenesis condition realising ST-SO10INLEP ST . We also show how this subset is not disrupted by the inclusion of flavour coupling but just slightly modified. For the time being, we will simply assume that the initial pre-existing asymmetry contribution is absent or in any case negligible.
The contribution to $\eta_{B}$ from leptogenesis, can be calculated as Buchmuller:2004nz
@@
This expression accounts for sphaleron conversion manton , kuzmin , klebnikov , Harvey:1990qw and photon dilution. The second numerical expression holds when the abundances $N_{X}$ are normalised in a portion of comoving volume such that the ultra-relativistic thermal equilibrium abundance of a RH neutrino $N_{N_{i}}^{\rm eq}(T\gg M_{i})=1$. Successful leptogenesis requires $\eta_{B}^{\rm lep}=\eta_{B}^{\rm exp}$, where $\eta_{B}^{\rm exp}$ is the measured value given in Eq. (24).
The final BLB-L asymmetry from leptogenesis can be calculated, neglecting flavour coupling, as the sum of the three contributions, one from each (charged lepton) flavour, explicitly vives , bounds , fuller , density :
@@
where $\Gamma_{I\alpha}=\Gamma(N_{I}\rightarrow\phi^{\dagger}\,l_{\alpha})$ and $\bar{\Gamma}_{I\alpha}=\Gamma(N_{I}\rightarrow\phi\,\bar{l}_{\alpha})$ are the zero temperature limit of the flavoured decay rates into $\alpha$ leptons and anti-leptons in the three-flavoured regime, $m_{\star}\equiv 16\,\pi^{5/2}\,\sqrt{g_{\star}}/(3\sqrt{5})\,(v^{2}/M_{\rm Pl})\simeq 1.08\,{\rm meV}$ is the equilibrium neutrino mass, $H(T)=\sqrt{g^{SM}_{\star}\,8\,\pi^{3}/90}\,T^{2}/M_{\rm P}$ is the expansion rate and $g_{\star}^{SM}=106.75$ is the number of ultra-relativistic degrees of freedom in the standard model. The quantities
@@
are the $N_{2}$-flavoured CPC\!P asymmetries, with $\Gamma_{2}\equiv\sum_{\alpha}\Gamma_{2\alpha}$ and $\overline{\Gamma}_{2}\equiv\sum_{\alpha}\,\overline{\Gamma}_{2\alpha}$. Finally, we defined $K_{2\tau_{2}^{\bot}}\equiv K_{2e}+K_{2\mu}$ and $\varepsilon_{2\tau_{2}^{\bot}}\equiv\varepsilon_{2e}+\varepsilon_{2\mu}$.
Using the bi-unitary parameterisation Eq. (12) for $m_{D}$, the flavoured decay parameters can be expressed as [^8] One can also express the decay parameters in terms of the orthogonal matrix: $K_{I\alpha}=\left|\sum_{j}\,\sqrt{m_{j}\over m_{\star}}\,U_{\nu\,\alpha j}\,\Omega_{jI}\right|^{2}\,,\hskip 14.22636pt\mbox{\rm and}\hskip 14.22636ptK_{I}=\sum_{I}\,{m_{j}\over m_{\star}}\,|\Omega_{jI}|^{2}\,.$ (30) However, in SO10INLEP, the bi-unitary parameterisation clearly provides the primary way to calculate all leptogenesis parameters, the $K_{I\alpha}$’s and the $\varepsilon_{2\alpha}$’s, since, as one can see from Eqs. (16) and (17), one gets directly RH neutrino masses $M_{I}$ and RH neutrino mixing matrix $U_{R}$ in terms of low energy neutrino parameters in $m_{\nu}$, neutrino Dirac masses $m_{Di}$ and parameters in $V_{L}$.
@@
The flavoured CPC\!P asymmetries can be calculated using crv
@@
with $\widetilde{m}_{2}\equiv(m_{D}^{{\dagger}}\,m_{D})_{22}/M_{2}$. However, the first term is important only when RH neutrino masses are quasi-degenerate, close to crossing level solutions. In our case such solutions are falling in region of parameter spaces that are now excluded by experiments [^9] For example, a crossing level solution where $\widetilde{m_{\nu}}^{-1}_{33}=0$ appears at $m_{1}\simeq 0.1\,{\rm eV}$ for the range of values of $\theta_{23}$ we are considering. In any case, in the code we use the full expression for $\varepsilon_{2\alpha}$, also including the contribution from the self-energy diagram to the CPC\!P asymmetries. and one can reliably use the approximate expression full
@@
Finally, for the efficiency factors at the production $\kappa(K_{2\alpha})$ we can use a simple analytic expression valid for initial thermal abundance flavourlep
$\kappa(K_{2\alpha})={2\over z_{B}(K_{2\alpha})\,K_{2\alpha}}\left(1-e^{-{K_{2\alpha}\,z_{B}(K_{2\alpha})\over 2}}\right)\,,\;\;z_{B}(K_{2\alpha})\simeq 2+4\,K_{2\alpha}^{0.13}\,e^{-{2.5\over K_{2\alpha}}}\,.$ (37)
Since all solutions are characterised by strong wash-out at the production (either $K_{2\tau}\gg 1$ or $K_{2\tau_{2}^{\bot}}\gg 1$ respectively for tauon and muon-dominated solutions), the final asymmetry does not depend on the initial $N_{2}$ abundance anyway. [^10] For $\tau_{B}$ solutions, that we introduce in the next section, one can find values of $K_{2\tau}$ as low as $K_{2\tau}\simeq 1$. However, solutions with $K_{2\tau}\lesssim 3$, where there is some dependence on the initial conditions and larger theoretical uncertainties, are realised only for values $m_{1}\gtrsim 30\,{\rm meV}$, now disfavoured by the cosmological upper bound Eq. (2).
We have now all the analytical expressions to calculate the final asymmetry in a fast and accurate way. Notice that in general the asymmetry within minimal leptogenesis with three RH neutrinos would depend on 18 quantities: $\varepsilon_{I\alpha},K_{I\alpha}$ ($I=1,2,3$ and $\alpha=e,\mu,\tau$). However, from Eq. (27) one can see that in SO10INLEP, since it realises a $N_{2}$-leptogenesis scenario, the asymmetry depends only on nine quantities, $\varepsilon_{2\alpha}\,K_{2\alpha},K_{1\alpha}\,(\alpha=e,\mu,\tau)$, while there is no dependence on $\varepsilon_{1\alpha},\varepsilon_{3\alpha},K_{3\alpha}(\alpha=e,\mu,\tau)$. These nine quantities can be expressed in terms of the seesaw parameters. As we have seen, using the bi-unitary parameterisation and the seesaw formula, one obtains Eqs. (16) and (17) for the three Majorana masses and the RH neutrino mixing matrix $U_{R}$ in terms of the six parameters in $V_{L}$, the nine parameters in the low energy neutrino matrix $m_{\nu}$ and the three Dirac neutrino masses $m_{Di}$. Therefore, in this way, the three RH neutrino masses and the six parameters in $U_{R}$ are traded off with the 9 low energy neutrino parameters. Moreover, the $\varepsilon_{2\alpha}$’s depend, approximately, only on $m_{D2}$ and the flavoured decay parameters do not depend on the Dirac neutrino masses $m_{D1},m_{D2}$ and $m_{D3}$. In this way, ultimately, one obtains for the final baryon-to-photon number ratio an expression of the form